OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

OpenStudy (mertsj):

You have 2 points so you need the slope: m = (6-0)/(2-4) = 6/-2 = -3

OpenStudy (anonymous):

y = 3(x + 4) y = –3(x – 4) y = 2x – 4 y = 4x – 2

OpenStudy (mertsj):

y-0=-3(x-4) And that is the equation in point-slope form.

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

Keke. You have it in slope-intercept form.

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

so its -3(x-4)